find the indicated outputs for f(x)=5x2-5x
f(0)=
f(-1)=
f(2)=

Answer by rapaljer(4671)

f(x)=5x^2-5x

f(0) means that x=0
f(0)=5*0^2-5*0=0

f(-1) means that x=-1
f(-1)=5(-1)^2-5(-1)
f(-1)=5*1+5=10

f(2) means that x=2
f(2)=5(2)^2-5(2)
f(-1)=5*4-10=10

Let me mention that math USUALLY looks harder than it really is, if it is taught right!! This is a concept that can really scare you, but in reality, it's not all that hard. Why don't you take a look at my explanation, and see if it is easier to understand than the one in your textbook. You can find my own website by clicking on my tutor name "rapaljer" anywhere in algebra.com. From here, you can see and click on my website.

From my Homepage, click on the link "Basic, Intermediate, and College Algebra: One Step at a time." Choose "Basic Algebra", and look in "Chapter 4" for "Section 4.09, Functional Notation". This topic is one of the SIMPLEST topics you can have in algebra, but it looks SO COMPLICATED. It's just the concept of substituting the given value in place of the x, then simplifying the result!! If you understand Signed Numbers and Order of Operations, then this should not be too hard to do. If you need help with Signed Numbers and Order of Operations, then you can an find my own explanation of these topics in Chapter 1 of my curriculum.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus